# The Mirrored Influence Hypothesis: Efficient Data Influence Estimation by Harnessing Forward Passes

### v1

@@ Abstract
In this paper, we introduce and explore the Mirrored Influence Hypothesis, highlighting a reciprocal nature of influence between training and test data. Specifically, it suggests that evaluating the influence of training data on test predictions can be reformulated as an equivalent, yet inverse problem: assessing how the predictions for training samples would be altered if the model were trained on specific test samples. Through both empirical and theoretical validations, we demonstrate the wide applicability of our hypothesis. Inspired by this, we introduce a new method for estimating the influence of training data, which requires calculating gradients for specific test samples, paired with a forward pass for each training point. This approach can capitalize on the common asymmetry in scenarios where the number of test samples under concurrent examination is much smaller than the scale of the training dataset, thus gaining a significant improvement in efficiency compared to existing approaches. We demonstrate the applicability of our method across a range of scenarios, including data attribution in diffusion models, data leakage detection, analysis of memorization, mislabeled data detection, and tracing behavior in language models. Our code will be made available at https://github.com/ruoxi-jia-group/Forward-INF.
@@ 4.2 Data Leakage Detection
Setup. In training data leakage evaluation, we use ResNet-18 (RN18) and ResNet-50 (RN50) classifiers [14], trained on CIFAR-10 [22] and CIFAR-100 [7], respectively. To simulate the case of data leakage, we set $D_{tst}\subset D_{trn}$. For baseline methods, we vary the key hyperparameters (i.e., the number of depths for Taylor expansion used by LISSA in approximating the Hessian for IF [2, 26] and the number of checkpoints for TracIn). We present the details, results of ImageNet100, and visualizations in Appendix E.
Result. As shown in Table 2, our approach is effective in identifying duplicated samples with 100% and 95% top-1 detection accuracy for both RN18 and RN50 classifiers. We observe that with the larger model RN50, the performance of IF deteriorates when detecting data leakage. This result is consistent with [5] that IFs are poor estimates of the impact of excluding a training point for neural networks. Although increasing the depth can help the algorithm in converging to a more accurate Hessian estimation, it results in an increased time requirement for each iteration. It is interesting to note that TracIn also exhibits poor performance in data leakage detection. Although it often identifies visually similar samples to the test example, it fails to accurately detect the ground-truth leaked sample. This discrepancy arises because, although the gradient of a duplicated sample may align with that of a test sample, numerous other training samples may align in a similar direction but with greater magnitude, yielding even higher scores than the duplicate itself. Additionally, TracIn in large-scale models often relies on last-layer gradient information  [3], which may lead to a suboptimal approximation to ground truth attribution. As shown in Table 8, Forward-INF achieves the best scores in terms of both efficacy and computational efficiency, while other approaches suffer from a significant computational bottleneck (benchmarked on NVIDIA GeForce RTX 2080 Ti), as well as difficulty in finding duplicated pairs.
@@ 4.3 Memorization Analysis
Results. In Figure 4, we compare the most influential training point as determined by our influence score calculations with the results from prior research. In each sub-figure, the first two panels display the test point and the most influential training point characterized by [11], while the third one presents the top-influence point retrieved by Forward-INF . Figure 4 shows that our approach can identify the same influential pairs as those in [11], but without retraining models. As mentioned in [11], the high-influential pairs (with an influence score $>0.4$) are near duplicated samples and benefit the most from memorization. Therefore, it leads to high memorization scores. Further qualitative results are provided in Appendix E.
@@ 5 Conclusion
Our contribution lies in the investigation of the Mirrored Influence Hypothesis. Expanding upon this hypothesis, we have developed a novel method to estimate train-to-test influence by solving the test-to-train influence problem. This approach involves evaluating the impact of incorporating a specific test set into the training set on the prediction of a training data source. Our method can be applied broadly and contribute meaningful insights across various settings. In particular, it outperforms traditional approaches that directly compute train-to-test influence by achieving an improved tradeoff between utility and computational efficiency.
Limitations & Future Work. The exploration of the Hypothesis unveils many avenues for future research. Notably, this paper does not incorporate heuristics to enhance Forward-INF. Nevertheless, it is anticipated that adopting strategic layer selection, akin to the approaches in TracIn, and integrating techniques to mitigate catastrophic forgetting [20] in the continual learning step, could potentially improve our method’s performance. Also, developing a theoretical framework to formally validate the Hypothesis presents an intriguing research direction.
@@ 6 Acknowledgment
We thank Hoang Anh Just and Himanshu Jahagirdar from the ReDS lab for their invaluable help in experiments and for their valuable discussion. RJ and the ReDS lab acknowledge support through grants from the Amazon-Virginia Tech Initiative for Efficient and Robust Machine Learning, the National Science Foundation under Grant No. IIS-2312794, NSF IIS-2313130, NSF OAC-2239622, and the CCI SWVA Research Engagement Award.
@@ Appendix A Related work
Figure 6: Correlation scores across different datasets under the noisy and near-noiseless settings. The first row shows the correlations between $\text{Inf}(D_{i}\rightarrow D_{\text{tst}})$ and $\text{Inf}(D_{i}\leftarrow D_{\text{tst}})$ under the near-noiseless setting, and the second row presents the correlation results under the noisy setting.
@@ Appendix D Hyperparameter Details [Section  4]
In this experiment, it is necessary to fine-tune a stable diffusion pre-trained model on a set of test samples. To fine-tune the pre-trained stable diffusion model, we follow the same setting as that from the previous work  [37]. In particular, we randomly pick fine-tuning exemplars from ImageNet [7] and further train the pre-trained model on the selected exemplars with 5 iterations, a learning rate of 1e-5, and a batch size of 4. After generating the synthesized samples, we perform gradient ascent on a set of synthesized samples with the same learning rate and 3 iterations.
@@ Appendix E Further Analysis and Details
Model/Dataset Metric IF-100 TracIn-1 TracIn-3 Forward-INF
ResNet-18 T-1 0.000 0.000 0.000 0.880
ImageNet-100 T-5 0.000 0.000 0.000 0.880
Figure 8: Performance comparison of ResNet-18 trained on the ImageNet-100 dataset. Here, TracIn 1 and TracIn 3 denote the number of checkpoints used for TracIn.
@@
Table 7: Top-1 detection accuracy for identifying the ground truth sample within candidate sets of varying sizes. Forward-INF shows 100% detection performance across different candidate set sizes.
@@
We further extend our experiment of data leakage detection on ImageNet-100 to validate the practicality as well as scalability. We observe a high detection rate (e.g., 88% top-1 detection rate) for the RN18 classifier trained on the ImageNet-100. As the model complexity increases, the detection rate slightly decreases, considering the previous results on CIFAR-10 and CIFAR-100. We hypothesize that the unlearning process affects more diverse neurons and their connections when larger and more complex datasets are used, leading to a drop in performance. However, this performance is still considered high, compared with our baselines such as IF and TracIN which completely fails even on this simple task.
@@
We additionally conduct a feature embedding analysis to visualize the gradient ascent dynamics in our method. In Figure  9, we present the feature space change according to the maximization steps. Interestingly, we observe that the duplicated sample (i.e., top-1) is strongly impacted by the gradient ascent process, experiencing a deviation from the class cluster during the process of unlearning. The figure in the bottom row shows examples that receive top-5 scores out of all training samples, indicating that our approach correctly identifies both the exact duplicated sample as well as relevant near duplicates. This observation supports our intuition that if we unlearn one specific point, we could expect that class or sample-related points would receive a high training loss change, thereby receiving a high Forward-INF score.
@@
Considering the importance of the selection of hyperparameters, we further vary the size of the validation set to verify the sensitivity of our design. As demonstrated in Table 8, our approach still provides 100% detection accuracy for our approach even with 25 validation samples, which supports that Forward-INF remains unaffected by changes in the size of the validation set.
@@
Table 8: Data leakage detection performance on ResNet-18 with different sizes of validation set for the hyperparameter selection. Forward-INF provides consistently high performance even for the small validation set size.

### v2

@@ Abstract
In this paper, we introduce and explore the Mirrored Influence Hypothesis, highlighting a reciprocal nature of influence between training and test data. Specifically, it suggests that evaluating the influence of training data on test predictions can be reformulated as an equivalent, yet inverse problem: assessing how the predictions for training samples would be altered if the model were trained on specific test samples. Through both empirical and theoretical validations, we demonstrate the wide applicability of our hypothesis. Inspired by this, we introduce a new method for estimating the influence of training data, which requires calculating gradients for specific test samples, paired with a forward pass for each training point. This approach can capitalize on the common asymmetry in scenarios where the number of test samples under concurrent examination is much smaller than the scale of the training dataset, thus gaining a significant improvement in efficiency compared to existing approaches. We demonstrate the applicability of our method across a range of scenarios, including data attribution in diffusion models, data leakage detection, analysis of memorization, mislabeled data detection, and tracing behavior in language models. Our code is available at https://github.com/ruoxi-jia-group/Forward-INF.
@@ 4.2 Data Leakage Detection
Setup. In training data leakage evaluation, we use ResNet-18 (RN18) and ResNet-50 (RN50) classifiers [14], trained on CIFAR-10 [22] and CIFAR-100 [7], respectively. To simulate the case of data leakage, we set $D_{tst}\subset D_{trn}$. For baseline methods, we vary the key hyperparameters (i.e., the number of depths for Taylor expansion used by LISSA in approximating the Hessian for IF [2, 26] and the number of checkpoints for TracIn). We present the details, and results of ImageNet100 in Appendix E.
Result. As shown in Table 2, our approach is effective in identifying duplicated samples with 100% and 95% top-1 detection accuracy for both RN18 and RN50 classifiers. We observe that with the larger model RN50, the performance of IF deteriorates when detecting data leakage. This result is consistent with [5] that IFs are poor estimates of the impact of excluding a training point for neural networks. Regarding the depth parameter, the algorithm’s convergence to $\alpha^{-1}(G+\lambda I)^{-1}v$ depends on the condition $\alpha(\tilde{G}+\lambda I)\preceq I$ being valid at every step  [grosse2023studying], which is rarely the case with large and complex models. Hence, errors could accumulate with each iteration, leading to the worst result for the IF-10000. Moreover using more iterations can result in a larger per-iteration cost. It is interesting to note that TracIn also exhibits poor performance in data leakage detection. Although it often identifies visually similar samples to the test example, it fails to accurately detect the ground-truth leaked sample as illustrated in Figure  12. This discrepancy arises because, although the gradient of a duplicated sample may align with that of a test sample, numerous other training samples may align in a similar direction but with greater magnitude, yielding even higher scores than the duplicate itself. Additionally, TracIn in large-scale models often relies on last-layer gradient information  [3], which may lead to a suboptimal approximation to ground truth attribution. As shown in Table 2, Forward-INF achieves the best scores in terms of both efficacy and computational efficiency, while other approaches suffer from a significant computational bottleneck (benchmarked on NVIDIA GeForce RTX 2080 Ti), as well as difficulty in finding duplicated pairs.
@@ 4.3 Memorization Analysis
Results. In Figure 4, we compare the most influential training point as determined by our influence score calculations with the results from prior research. In each sub-figure, the first two panels display the test point and the most influential training point characterized by [11], while the third one presents the top-influence point retrieved by Forward-INF . Figure 4 shows that our approach can identify the same influential pairs as those in [11] but without retraining models. As mentioned in [11], the high-influential pairs (with an influence score $>0.4$) are near duplicated samples and benefit the most from memorization. Therefore, it leads to high memorization scores. Further qualitative results are provided in Appendix E.
@@ 5 Conclusion
Our contribution lies in the investigation of the Mirrored Influence Hypothesis. Expanding upon this hypothesis, we have developed a novel method to estimate train-to-test influence by solving the test-to-train influence problem. This approach involves evaluating the impact of incorporating a specific test set into the training set on the prediction of a training data source. Our method can be applied broadly and contribute meaningful insights across various settings. In particular, it outperforms traditional approaches that directly compute train-to-test influence by achieving an improved tradeoff between utility and efficiency.
Limitations & Future Work. The exploration of the Hypothesis unveils many avenues for future research. Although this paper excludes heuristic enhancements to Forward-INF , it is expected that strategic layer selection, and incorporating strategies to counter catastrophic forgetting [20], could enhance our method’s performance. Formulating a theoretical framework to formally validate the Hypothesis constitutes an intriguing direction for future studies.
@@ 6 Acknowledgment
We thank Hoang Anh Just and Himanshu Jahagirdar from the ReDS lab for their invaluable help in experiments and discussion. RJ and the ReDS lab acknowledge support through grants from the Amazon-Virginia Tech Initiative for Efficient and Robust Machine Learning, the National Science Foundation under Grant No. IIS-2312794, NSF IIS-2313130, NSF OAC-2239622, and the CCI SWVA Research Engagement Award.
@@ Appendix A Related work
Figure 6: Correlation scores across different datasets under the noisy and near-noiseless settings. The first row shows the correlations between $\text{Inf}(D_{i}\rightarrow D_{\text{tst}})$ and $\text{Inf}(D_{i}\leftarrow D_{\text{tst}})$ under the near-noiseless setting, and the second row presents the correlation results under the noisy setting. We note that the correlation scores for CIFAR-10 under the noisy setting can be increased to 0.8971, and 0.9551 respectively by changing the number of epochs.
@@ Appendix D Hyperparameter Details [Section  4]
In this experiment, it is necessary to fine-tune a stable diffusion pre-trained model on a set of test samples. To fine-tune the pre-trained stable diffusion model, we follow the same setting as that from the previous work  [37]. In particular, we randomly pick fine-tuning exemplars from ImageNet [7] and further train the pre-trained stable diffusion model on the selected exemplars with 5 iterations, a learning rate of 1e-5, and a batch size of 4. After generating the synthesized samples, we perform gradient ascent on a set of synthesized samples with the same learning rate and 3 iterations.
@@ Appendix E Further Analysis and Details
Model/Dataset Metric IF-100 Forward-INF
ResNet-18 T-1 0.000 0.880
ImageNet-100 T-5 0.000 0.880
Table 6: Data Leakage detection performance comparison of ResNet-18 trained on the ImageNet-100 dataset.
@@
Table 8: Top-1 detection accuracy for identifying the ground truth influential sample within candidate sets of varying sizes for the stable diffusion model. Forward-INF shows 100% detection performance across different candidate set sizes.
Figure 12: TracIn and IF results on data leakage experiments.The right side of the figure illustrates that even though TracIn can often identify visually similar samples to the test example, it fails to detect the ground-truth leaked sample. The left side of the figure showcases the results using different depths and scales to empirically demonstrate why IF-10000 falls short.
@@
We further extend our experiment of data leakage detection on ImageNet-100 to validate the practicality as well as scalability. We observe a high detection rate (e.g., 88% top-1 detection rate) for the RN18 classifier trained on the ImageNet-100. As the model complexity increases, the detection rate slightly decreases, considering the previous results on CIFAR-10 and CIFAR-100. We hypothesize that the unlearning process affects more diverse neurons and their connections when larger and more complex datasets are used, leading to a drop in performance. However, this performance is still considered high, compared with our baseline such as IF which completely fails as shown in Table  6.
@@
We additionally conduct a feature embedding analysis to visualize the gradient ascent dynamics in our method. In Figure  8, we present the feature space change according to the maximization steps. Interestingly, we observe that the duplicated sample (i.e., top-1) is strongly impacted by the gradient ascent process, experiencing a deviation from the class cluster during the process of unlearning. The figure in the bottom row shows examples that receive top-5 scores out of all training samples, indicating that our approach correctly identifies both the exact duplicated sample as well as relevant near duplicates. This observation supports our intuition that if we unlearn one specific point, we could expect that class or sample-related points would receive a high training loss change, thereby receiving a high Forward-INF score.
@@
Considering that the unlearning process may involve sensitivity to hyperparameters (e.g., learning rate, and the number of iterations), our method initially selects the hyperparameter based on a holdout validation set. Consequently, we also present the results concerning the sensitivity of the validation set size. As illustrated in Table  9, Forward-INF still achieves 100% detection accuracy for our approach even with only 25 validation samples, indicating that our approach exhibits low sensitivity to changes in the size of the validation set.
@@
Table 9: Data leakage detection performance using ResNet-18 across various validation set sizes for the sensitivity analysis of hyperparameter selection. Forward-INF consistently delivers high performance, even with small validation set sizes.